OpenStudy (anonymous):

tanθ=-√7/3 and θ is in quad IV. Find secθ

OpenStudy (anonymous):

\[\sec^2\theta= 1+\tan^2\theta\]

OpenStudy (anonymous):

sec is positive in fourth quad, just put the value and take the sqrt

OpenStudy (anonymous):

4/3?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Another slant: \[\text{Sec}\left[\text{ArcTan}\left[-\frac{\sqrt{7}}{3}\right]\right]=\frac{4}{3} \]